I am trying to solve this story problem and I am not getting the right answer no matter what I do. Just need some help.
Soybean meal is 18% protein and cornmeal is 9% protein. How many pounds of each should be mixed together in order to get 360-lb mixture that is 13% protein?
Thanks :o)

Answer by rfer(16322)

.18x+.09(360-x)=.13(360)
.18x+32.4-.09x=46.8
.09x=14.4
x=160 lbs of soybean meal
360-160=200 lbs cornmeal